# Doping of a Borexino-like Liquid Scintillator with Tellurium-Diols

### v2

@@ Acknowledgments
This work benefited substantially from the support with seed funding from the Cluster of Excellence ORIGINS (DFG, German Research Foundation under Germany’s Excellence Strategy – EXC-2094 – 390783311).
This work benefited from the support by the Detector Laboratory of the Cluster of Excellence PRISMA+. We are very grateful to the staff of the Detector Division around Dr. Quirin Weitzel. We would like to express our warmest thanks to Dr. Mihail Mondeshki (Chemistry Department, JGU Mainz) and his team for their support with chemical expertise, their unbiased interdisciplinary collaboration, and numerous discussions.
We are grateful to the University of Milan and to INFN of Milan which host the fluorescence time profile setup in the Liquid Scintillator laboratory. We would also thank Dr. Barbara Cacciania and Fatima Horuia for the help for this work.

### v3

@@ Acknowledgments
We gratefully acknowledge seed funding from the Munich Cluster of Excellence ORIGINS (DFG, German Research Foundation under Germany’s Excellence Strategy – EXC-2094 – 390783311) and the support of the Detector Laboratory of the Mainz Cluster of Excellence PRISMA+.
Special thanks go to the staff of the Detector Division, in particular Dr. Quirin Weitzel, for their practical assistance. We also warmly thank Dr. Mihail Mondeshki (Chemistry Department, JGU Mainz) and his team for their chemical expertise, constructive interdisciplinary collaboration, and numerous insightful discussions. We would like to thank the chemotechnical employee of the Mainz Institute for Physics, Joachim Strübig, and the glass apparatus manufacturer Rainer Jera.
We are grateful to the University of Milan and to INFN of Milan which host the fluorescence time profile setup in the Liquid Scintillator laboratory. We would also thank Dr. Barbara Caccianiga and Fatima Horuia for their help for this work.